solve using the elimination method.
8x-3y=-6
5x+6y=75

Answer by dkala99(12)

8x-3y=-6 ----> 1
5x+6y=75 ----> 2
Mutiply equation 1 by 2 on both sides
16x-6y=-12 ----> 3
Add equation 2 and 3
5x+6y=75 ----> 2
16x-6y=-12 ----> 3
------------------
21x = 63
x = 63/21 = 3
x = 3
Apply x=3 in equation 1
8*3 - 3y = -6
24 - 3y = -6
-3y = -6-24 = -30
y = -30/-3
y = 10
x = 3 ; y = 10